OpenStudy (anonymous):

Slope of tangent f(x) at x=a is fprime(a). Can you proof tangent line equation ? y=f(a)+fprime(a)(x-a)

OpenStudy (anonymous):

hope you understan my handwriting...

OpenStudy (anonymous):

f(a+h) - f(a) slope of the tangent m=(lim as h->0) ---------- h

OpenStudy (anonymous):

therefore y=f(a)+f '(a)(x-a)

OpenStudy (anonymous):

hi tech is this the correct prob? y=f(a)+f '(a)(x-a)